# thanks: Associate Member of JETSCAPEthanks: Associate Member of JETSCAPEthanks: Associate Member of JETSCAPEthanks: Associate Member of JETSCAPE

### v1

@@ Abstract
We employ the SMASH transport model to provide event-by-event initial conditions for the energy-momentum tensor and conserved charge currents in hydrodynamic simulations of relativistic heavy-ion collisions. We study the fluctuations and dynamical evolution of three conserved charge currents (net baryon, net electric charges, and net strangeness) with a 4D lattice-QCD-based equation of state, NEOS-4D, in the hydrodynamic phase. Out-of-equilibrium corrections at the particlization are generalized to finite densities to ensure the conservation of energy, momentum, and the three types of charges. These theoretical developments are integrated within X-SCAPE as a unified framework for studying the nuclear matter properties in the Beam Energy Scan program.
@@ II Model Description
In this work, we employ the microscopic non-equilibrium hadronic transport model SMASH [35] as the initial conditions for relativistic viscous hydrodynamics, followed by particlization using the Cooper-Frye procedure [57] when the energy densities of fluid cells drop below a switching energy density. The produced hadrons are then fed back to SMASH and evolved until their kinetic freeze-out and final strong decays are performed. This setup is similar to other hybrid approaches with hadronic transport initial conditions, like the SMASH-vHLLE-Hybrid [42, 58] or setups using UrQMD in the initial state [37, 24, 40, 59].
@@ II.1 Hadronic Transport
The hadronic transport model SMASH is designed to evolve a hadronic out-of-equilibrium system, as it can be found in the initial and final state of heavy-ion collisions, in phase space [35, 60]. SMASH is an effective solver for the system of relativistic Boltzmann equations
$p^{\mu}\frac{\partial}{\partial x^{\mu}}f_{i}(t,\mathbf{x},\mathbf{p})+m_{i}F^{\alpha}\frac{\partial}{\partial p^{\alpha}}f_{i}(t,\mathbf{x},\mathbf{p})=\mathcal{C}^{i}\left[f_{i}(t,\mathbf{x},\mathbf{p})\right],$ (1)
where $i$ labels the hadron species. The collision integral $\mathcal{C}^{i}\left[f_{i}(t,\mathbf{x},\mathbf{p})\right]$, which is a functional of the one-particle distribution function, is modeled by formations, decays, and elastic scatterings of hadronic resonances. $p^{\mu}$ denotes the particle’s four-momentum, $m_{i}$ the mass of particles of the hadron species $i$ and $F^{\mu}$ is an external force. In this work, we employ only the cascade mode of the transport code, ensuring that the force term in the Boltzmann equation does not impact the evolution in our setup. SMASH incorporates all hadronic degrees of freedom up to a mass of $2.35\;\mathrm{GeV}$ listed in the PDG table [61]. The high-energy hadronic interactions in SMASH are treated with string formation and fragmentation using Pythia 8 [62, 63]. Interactions at lower energies are implemented using resonances with vacuum Breit-Wigner spectral functions. These have mass-dependent widths implemented with the Manley-Saleski ansatz [64]. The collisions between hadrons use a geometric covariant collision criterion [65].
@@ II.1.1 Initial Conditions
The colliding nuclei in SMASH are sampled from a Woods–Saxon distribution, shifted by the impact parameter, and propagated along the $z$-direction. For the initialization of relativistic hydrodynamics, particles are extracted from SMASH on a constant proper time hypersurface $\tau_{0}$, defined by the nuclear passing time [24]:
@@
where $R_{\rm p/t}$ are the projectile and target radii, $\sqrt{s_{\rm NN}}$ the center-of-mass energy, and $m_{\rm N}$ the nucleon mass. Ref. [42] demonstrated that this removal, which leaves holes in the hadronic medium, does not affect final-state observables. At high collision energies, for example $\sqrt{s_{\rm NN}}=200\;\mathrm{GeV}$, we impose a lower bound of $\tau_{0}=0.5\;\mathrm{fm}/c$. [^1] For comparison, the values $\sqrt{s_{\rm NN}}=19.6\,\mathrm{GeV}$ and $7.7\,\mathrm{GeV}$ correspond to initialization times of $\tau_{0}=1.23\,\mathrm{fm}/c$ and $3.2\,\mathrm{fm}/c$, respectively, based on the nucleon mass $m_{\mathrm{N}}=0.938\,\mathrm{GeV}$ and gold radius $R_{\rm Au}=6.37\,\mathrm{fm}$ as implemented in SMASH.
To prevent nuclei without potentials from disintegrating before the collision, the initial Fermi motion in the transverse plane is frozen. It is reactivated as soon as the first nucleon–nucleon scattering occurs.
@@ II.2 Hydrodynamics
The four-position of the hadron is indicated by $x^{\mu}_{\rm h}$. In this work, we employ a covariant smearing kernel and compare its results with those from a regular Gaussian smearing kernel. The covariant smearing kernel in Cartesian coordinates was employed in Refs. [70, 41]:
@@
The hydrodynamic equation of motion in Eqns. (3) and (4) are solved together with an equation of state neos-4D based on lattice QCD and hadron resonance gas at finite densities [52, 56]. The complete 4-dimensional dependence for pressure on local energy density and three types of conserved charges, $P(e,n_{B},n_{Q},n_{S})$, enables the independent propagation of net baryon, electric charge, and strangeness inside the QCD fluid. We consider viscous effects, namely shear and bulk viscosity, in fluid dynamic evolution. The shear stress tensor and bulk viscous pressure are evolved with the Denicol-Niemi-Molnar-Rischke (DNMR) theory [72, 73]. The specific shear and bulk viscosity are parametrized as
@@ III X-SCAPE Implementation
The simulation begins by initializing the nuclei in SMASH according to a Woods–Saxon distribution, separated by the chosen impact parameter, which can be either fixed or randomly sampled within a range. The time origin $t=0$ is then defined dynamically as the moment of the first nucleon–nucleon collision. During the pre-equilibrium stage, SMASH propagates the produced hadrons in Minkowski coordinates, while the BDM module continuously extracts particles crossing the iso-$\tau_{0}=0.5\,\mathrm{fm}/c$ hypersurface. To ensure accurate determination of hadron crossings on this surface, the pre-equilibrium hadronic transport evolution is performed with a very small timestep of $\Delta t=0.002\;\mathrm{fm}/c$. Hadrons in the extracted hadron list are back-propagated using free-streaming, ensuring they are exactly on the iso-$\tau_{0}$ hypersurface when the source terms are created. All participant hadrons are then given to a hadron liquefier module that smears the source terms according to the procedure described in Sec. II.2. At the same time, the spectator hadrons are stored for possible use in the afterburner phase. [^2] In this work, we do not include the spectator hadrons in the afterburner phase.
@@ V Conclusions
To capture all the finite density dependence in the full-fledged bulk dynamics simulations, we couple the SMASH initial conditions with relativistic hydrodynamics using a 4D lattice-QCD-based equation of state, in which the local pressure depends on $(e,n_{B},n_{Q},n_{S})$. We further generalize the out-of-equilibrium corrections with all three types of conserved charges at finite densities in the Cooper-Frye particlization procedure. They are essential to ensure all the conserved macroscopic quantities are smoothly switched from fluid cells to individual hadrons. Using this new hybrid model, we conduct a pilot study of the (3+1)D dynamical evolution of all three conserved charge densities, event by event. We investigate the dynamical evolution of the system’s entropy density and the three types of conserved charges during hydrodynamic evolution at several collision energies. We observe that the local charge fluctuations remain substantial for net electric charges and net strangeness across all collision energies. The full covariant smear kernel exhibits large pressure gradients, resulting in strong hydrodynamic radial flow and flattened particle spectra.
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant number OAC-2004571 (CSSI:X-SCAPE). It was also supported under PHY-1516590 and PHY-1812431 (R.J.F., M.Ko., C.P. and A.S.); it was supported in part by the US Department of Energy, Office of Science, Office of Nuclear Physics under grant numbers DE-AC02-05CH11231 (X.-N.W. and W.Z.), DE-AC52-07NA27344 (D.A.H. and R.A.S.), DE-SC0013460 (A.M., I.S., C.Si., R.Da. and R.Do.), DE-SC0021969 (C.Sh., G.P., and W.Z.), DE-SC0024232 (C.Sh. and H.R.), DE-SC0012704 (B.S.), DE-FG02-92ER40713 (J.H.P.), DE-FG02-05ER41367 (C.Si and S.A.B.), DE-SC0024660 (R.K.E), DE-SC0024347 (J.-F.P. and M.S.). The work was also supported in part by the National Science Foundation of China (NSFC) under grant numbers 11935007, 11861131009 and 11890714 (X.-N.W.), by the Natural Sciences and Engineering Research Council of Canada (C.G., S.J., and G.V.), by the University of Regina President’s Tri-Agency Grant Support Program (G.V.), by the Canada Research Chair program (G.V. and A.K.) reference number CRC-2022-00146, by JSPS KAKENHI Grant Numbers 22K14041 and 25K07303 (Y.T.), JP19K14722 and JP24K07030 (A.Mo.), and by the São Paulo Research Foundation (FAPESP) under projects 2016/24029-6, 2017/05685-2 and 2018/24720-6 (M.L.). C.Sh., J.-F.P., and R.K.E. acknowledge a DOE Office of Science Early Career Award. I. S. was funded as part of the European Research Council project ERC-2018-ADG-835105 YoctoLHC, and as a part of the Center of Excellence in Quark Matter of the Academy of Finland (project 346325). Calculations for this work used the Wayne State Grid, generously supported by the Office of the Vice President of Research (OVPR) at Wayne State University.

### v2

@@ Abstract
We employ the SMASH transport model to provide event-by-event initial conditions for the energy-momentum tensor and conserved charge currents in hydrodynamic simulations of relativistic heavy-ion collisions. We study the fluctuations and dynamical evolution of three conserved charge currents (net baryon, net electric charges, and net strangeness) with a 4D lattice-QCD-based equation of state, NEOS-4D, in the hydrodynamic phase. Out-of-equilibrium corrections at the particlization are generalized to finite densities to ensure the conservation of energy, momentum, and the three types of charges. These theoretical developments are integrated within the X-SCAPE code as a unified framework for studying the nuclear matter properties in the Beam Energy Scan program.
@@ II Model Description
In this work, we employ the microscopic non-equilibrium hadronic transport model SMASH [35] (version 3.1) as the initial conditions for relativistic viscous hydrodynamics, followed by particlization using the Cooper-Frye procedure [57] when the energy densities of fluid cells drop below a switching energy density. The produced hadrons are then fed back to SMASH and evolved until their kinetic freeze-out and final strong decays are performed. This setup is similar to other hybrid approaches with hadronic transport initial conditions, like the SMASH-vHLLE-Hybrid [42, 58] or setups using UrQMD in the initial state [37, 24, 40, 59].
@@ II.1 Hadronic Transport
The hadronic transport model SMASH is designed to evolve a hadronic out-of-equilibrium system, as it can be found in the initial and final state of heavy-ion collisions, in phase space [35, 60]. In this work, we run SMASH with physical particles in the quantum molecular dynamics (QMD) mode without mean field and test particle method. SMASH incorporates all hadronic degrees of freedom up to a mass of $2.35\;\mathrm{GeV}$ listed in the PDG table [61]. The high-energy hadronic interactions in SMASH are treated with string formation and fragmentation using Pythia 8 [62, 63]. Interactions at lower energies are implemented using resonances with vacuum Breit-Wigner spectral functions. These have mass-dependent widths implemented with the Manley-Saleski ansatz [64]. The collisions between hadrons use a geometric covariant collision criterion [65] including $1\rightarrow 2$, $2\rightarrow 1$, and $2\rightarrow 2$ processes.
@@ II.1.1 Initial Conditions
The colliding nuclei in SMASH are sampled from a Woods–Saxon distribution, shifted by the impact parameter, and propagated along the $z$-direction. To prevent nuclei without potentials from disintegrating before the collision, the initial Fermi motion in the transverse plane is frozen. It is reactivated as soon as the first nucleon–nucleon scattering occurs.
For the initialization of relativistic hydrodynamics, particles are extracted from SMASH on a constant proper time hypersurface $\tau_{0}$, defined by the nuclear passing time [24]:
@@
where $R_{\rm p/t}$ are the projectile and target radii, $\sqrt{s_{\rm NN}}$ the center-of-mass energy, and $m_{\rm N}$ the nucleon mass. At high collision energies, for example $\sqrt{s_{\rm NN}}=200\;\mathrm{GeV}$, we impose a lower bound of $\tau_{0}=0.5\;\mathrm{fm}/c$ similar to Ref. [42]. [^1] For comparison, the values $\sqrt{s_{\rm NN}}=19.6\,\mathrm{GeV}$ and $7.7\,\mathrm{GeV}$ correspond to initialization times of $\tau_{0}=1.23\,\mathrm{fm}/c$ and $3.2\,\mathrm{fm}/c$, respectively, based on the nucleon mass $m_{\mathrm{N}}=0.938\,\mathrm{GeV}$ and gold radius $R_{\rm Au}=6.37\,\mathrm{fm}$ as implemented in SMASH.
At high energies, most of the hadrons are produced by string fragmentation [66]. SMASH assumes that strings are decayed into hadrons instantaneously through Pythia. Therefore, this setup does not include overlapping and interacting strings [67]. Using SMASH as an initial condition effectively sets the hadron formation time to $\tau_{0}$. The physical process of string fragmentation requires finite time, which could be larger than $\tau_{0}$. Our approach assumes a fast chemical equilibration from a gluon-dominated phase. We use only the space-time information of the hadrons’ energies and the quantum numbers of the conserved charges to initialize the hydrodynamic fields.
@@ II.2 Hydrodynamics
The four-position of the hadron is indicated by $x^{\mu}_{\rm h}$. Since we are dealing with hadrons, there is no pre-equilibrium color field contribution to the hydrodynamic source terms. In this work, we employ a covariant smearing kernel and compare its results with those from a regular Gaussian smearing kernel. The covariant smearing kernel in Cartesian coordinates was employed in Refs. [72, 41]:
@@
The hydrodynamic equation of motion in Eqns. (2) and (3) are solved together with an equation of state neos-4D based on lattice QCD and hadron resonance gas at finite densities [52, 56]. This approach assumes that the system reaches chemical equilibrium at $\tau_{0}$. We leave the exploration of dynamical chemical equilibration for future work. The complete 4-dimensional dependence for pressure on local energy density and three types of conserved charges, $P(e,n_{B},n_{Q},n_{S})$, enables the independent propagation of net baryon, electric charge, and strangeness inside the QCD fluid. We consider viscous effects, namely shear and bulk viscosity, in fluid dynamic evolution. The shear stress tensor and bulk viscous pressure are evolved with the Denicol-Niemi-Molnar-Rischke (DNMR) theory [74, 75]. The specific shear and bulk viscosity are parametrized as
@@ II.3 Particlization
We use the $(e,n_{B},n_{Q},n_{S})$ in individual hydrodynamic cells and convert them to $(T,\mu_{B},\mu_{Q},\mu_{S})$ using a hadron resonance gas (HRG) EOS including the hadronic content of the SMASH transport code. [^2] The HRG does not include phase shift corrections from hadronic interactions, but the resonance states effectively take into account some of them [80]. This procedure ensures the conservation of energy-momentum and conserved charges during the particlization stage. The official neos-4D matches to an HRG with a slightly different hadronic content, leading to a less than 5% difference in the conversion. Ref. [81] shows that the discrepancy is negligible in the final-state observables.
@@ III X-SCAPE Implementation
The simulation begins by initializing the nuclei in SMASH according to a Woods–Saxon distribution, separated by the chosen impact parameter, which can be either fixed or randomly sampled within a range. The time origin $t=0$ is then defined dynamically as the moment of the first nucleon–nucleon collision. During the pre-equilibrium stage, SMASH propagates the produced hadrons in Minkowski coordinates, while the BDM module continuously extracts particles crossing the iso-$\tau_{0}=0.5\,\mathrm{fm}/c$ hypersurface. Ref. [42] demonstrated that this removal, which leaves holes in the hadronic medium, does not affect final-state observables. To ensure accurate determination of hadron crossings on this surface, the pre-equilibrium hadronic transport evolution is performed with a very small timestep of $\Delta t=0.002\;\mathrm{fm}/c$. Hadrons in the extracted hadron list are back-propagated using free-streaming, ensuring they are exactly on the iso-$\tau_{0}$ hypersurface when the source terms are created. All participant hadrons are then given to a hadron liquefier module that smears the source terms according to the procedure described in Sec. II.2. At the same time, the spectator hadrons are stored for possible use in the afterburner phase. [^3] In this work, we do not include the spectator hadrons in the afterburner phase.
@@ V Conclusions
To capture all the finite density dependence in the full-fledged bulk dynamics simulations, we couple the SMASH initial conditions with relativistic hydrodynamics using a 4D lattice-QCD-based equation of state, in which the local pressure depends on $(e,n_{B},n_{Q},n_{S})$. We further generalize the out-of-equilibrium corrections with all three types of conserved charges at finite densities in the Cooper-Frye particlization procedure. They are essential to ensure all the conserved macroscopic quantities are smoothly switched from fluid cells to individual hadrons. Using this new hybrid model, we conduct a pilot study of the (3+1)D dynamical evolution of all three conserved charge densities in the hydrodynamic phase and its subsequent impact on final-state observables. We investigate the dynamical evolution of the system’s entropy density and the three types of conserved charges during hydrodynamic evolution at several collision energies. We observe that the local charge fluctuations remain substantial for net electric charges and net strangeness across all collision energies. The full covariant smear kernel exhibits large pressure gradients, resulting in strong hydrodynamic radial flow and flattened particle spectra.
@@
The framework code and parameter input files used to reproduce the results of this work are available in Refs. [90, 107].
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant numbers OAC-2004571 (CSSI:X-SCAPE) and OAC-2514008 (CSSI:C-SCAPE). It was also supported under PHY-2111568 and PHY-2413003 (R.J.F., M.Ko., C.P. and A.S.); it was supported in part by the US Department of Energy, Office of Science, Office of Nuclear Physics under grant numbers DE-AC02-05CH11231 (X.-N.W. and W.Z.), DE-AC52-07NA27344 (D.A.H. and R.A.S.), DE-SC0013460 (A.M., I.S., C.Si., R.Da. and R.Do.), DE-SC0021969 (C.Sh., G.P., and W.Z.), DE-SC0024232 (C.Sh. and H.R.), DE-SC0012704 (B.S.), DE-FG02-92ER40713 (J.H.P.), DE-FG02-05ER41367 (C.Si and S.A.B.), DE-SC0024660 (R.K.E), DE-SC0024347 (J.-F.P. and M.S.). The work was also supported in part by the National Science Foundation of China (NSFC) under grant numbers 11935007, 11861131009 and 11890714 (X.-N.W.), by the Natural Sciences and Engineering Research Council of Canada (C.G., S.J., and G.V.), by the University of Regina President’s Tri-Agency Grant Support Program (G.V.), by the Canada Research Chair program (G.V. and A.K.) reference number CRC-2022-00146, by JSPS KAKENHI Grant Numbers 22K14041 and 25K07303 (Y.T.), JP19K14722 and JP24K07030 (A.Mo.), and by the São Paulo Research Foundation (FAPESP) under projects 2016/24029-6, 2017/05685-2 and 2018/24720-6 (M.L.). C.Sh., J.-F.P., and R.K.E. acknowledge a DOE Office of Science Early Career Award. I. S. was funded as part of the European Research Council project ERC-2018-ADG-835105 YoctoLHC, and as a part of the Center of Excellence in Quark Matter of the Academy of Finland (project 346325). Calculations for this work used the Wayne State Grid, generously supported by the Office of the Vice President of Research (OVPR) at Wayne State University.